# TwinMarket: A Scalable Behavioral and Social Simulation for Financial Markets

### v2

@@ B.1 SSE 50 and 10 Aggregated Index
The SSE 50 index is a benchmark stock market index comprising the 50 largest and most liquid A-share stocks listed on the Shanghai Stock Exchange. It is widely used to track the performance of the blue-chip segment of the Chinese stock market.
The 10 aggregated indexes, on the other hand, are designed to represent different sectors or industries within the SSE 50. Each aggregated index consists of a subset of SSE 50 constituent stocks, grouped based on their industry classifications. The composition and weighting of each index are detailed below:
  • TLEI (Transportation and Logistics Index): This index comprises 2 constituent stocks, including COSCO SHIPPING Holdings (SH601919, weight 52.34%) and China State Shipbuilding Corporation (SH600150, weight 47.66%).
  • MEI (Manufacturing Index): This index comprises 8 constituent stocks, including LONGi Green Energy Technology (SH601012, weight 16.51%), Haier Smart Home (SH600690, weight 15.78%), SANY Heavy Industry (SH600031, weight 15.29%), NARI Technology (SH600406, weight 14.6%), SAIC Motor (SH600104, weight 11.98%), Tongwei (SH600438, weight 10.92%), TBEA (SH600089, weight 10.12%), and Great Wall Motor (SH601633, weight 4.8%).
  • CPEI (Chemical and Pharmaceutical Index): This index comprises 3 constituent stocks, including Jiangsu Hengrui Pharmaceuticals (SH600276, weight 45.93%), Wanhua Chemical Group (SH600309, weight 28.39%), and WuXi AppTec (SH603259, weight 25.68%).
  • IEEI (Infrastructure and Engineering Index): This index comprises 3 constituent stocks, including China State Construction Engineering (SH601668, weight 52.25%), China Railway Group (SH601390, weight 27.64%), and Power Construction Corporation of China (SH601669, weight 20.11%).
  • REEI (Real Estate Index): This index comprises 1 constituent stock, including Poly Developments and Holdings (SH600048, weight 100.0%).
@@
  • TSEI (Tourism and Service Index): This index comprises 1 constituent stock, including China Tourism Group Duty Free (SH601888, weight 100.0%).
  • CGEI (Consumer Goods Index): This index comprises 5 constituent stocks, including Kweichow Moutai (SH600519, weight 69.18%), Inner Mongolia Yili Industrial Group (SH600887, weight 13.13%), Shanxi Xinghuacun Fen Wine Factory (SH600809, weight 7.17%), Foshan Haitian Flavouring and Food (SH603288, weight 5.44%), and Zhangzhou Pientzehuang Pharmaceutical (SH600436, weight 5.08%).
  • TTEI (Technology and Telecommunications Index): This index comprises 10 constituent stocks, including Semiconductor Manufacturing International Corporation (SH688981, weight 18.07%), Hygon Information Technology (SH688041, weight 11.89%), China Telecom (SH601728, weight 10.21%), China Unicom (SH600050, weight 9.98%), Advanced Micro-Fabrication Equipment Inc. China (SH688012, weight 9.79%), China Mobile (SH600941, weight 9.76%), China National Nuclear Power (SH601985, weight 9.05%), Will Semiconductor (SH603501, weight 8.53%), Kingsoft Office Software (SH688111, weight 6.92%), and GigaDevice Semiconductor (SH603986, weight 5.81%).
  • EREI (Energy and Resources Index): This index comprises 6 constituent stocks, including China Yangtze Power (SH600900, weight 33.45%), Zijin Mining Group (SH601899, weight 25.88%), China Shenhua Energy (SH601088, weight 13.19%), Sinopec (SH600028, weight 9.29%), PetroChina (SH601857, weight 9.12%), and Shaanxi Coal Industry (SH601225, weight 9.07%).
  • FSEI (Financial Services Index): This index comprises 11 constituent stocks, including Ping An Insurance (SH601318, weight 22.86%), China Merchants Bank (SH600036, weight 17.94%), CITIC Securities (SH600030, weight 11.97%), Industrial Bank (SH601166, weight 10.47%), Industrial and Commercial Bank of China (SH601398, weight 8.6%), Bank of Communications (SH601328, weight 8.04%), Agricultural Bank of China (SH601288, weight 6.12%), China Pacific Insurance (SH601601, weight 4.63%), Bank of China (SH601988, weight 4.21%), China Life Insurance (SH601628, weight 2.8%), and Postal Savings Bank of China (SH601658, weight 2.35%).
B.2 Real Stock Data Variables
The trading data is collected on a daily basis and includes the following variables for each stock:
  • stock_id: The unique identifier code for each stock.
  • close_price: The closing price of the stock on the trading day.
  • pre_close: The closing price of the stock on the previous trading day.
  • change: The change in price of the stock from the previous trading day’s close to the current day’s close.
  • pct_chg: The percentage change in the stock’s price from the previous trading day’s close to the current day’s close.
  • pe_ttm: The price-to-earnings ratio calculated using trailing twelve months (TTM) earnings.
  • pb: The price-to-book ratio.
  • ps_ttm: The price-to-sales ratio calculated using trailing twelve months (TTM) sales.
  • dv_ttm: The dividend yield calculated using trailing twelve months (TTM) dividends.
  • vol: The trading volume of the stock on the trading day.
  • vol_5: The 5-day average trading volume.
  • vol_10: The 10-day average trading volume.
  • vol_30: The 30-day average trading volume.
  • ma_hfq_5: The 5-day moving average of the stock’s closing price, using the backward-adjusted method to account for stock splits and dividends.
  • ma_hfq_10: The 10-day moving average of the stock’s closing price, using the backward-adjusted method.
  • ma_hfq_30: The 30-day moving average of the stock’s closing price, using the backward-adjusted method.
  • elg_amount_net: The net inflow of large orders (if available). This represents the difference between buy and sell orders of a certain size that are considered to be institutional orders.
  • date: The trading date associated with the record.
In addition to the trading data, the dataset includes basic information about each company. These variables are generally static but may be updated periodically:
  • reg_capital: The registered capital of the company (in RMB).
  • setup_date: The date when the company was established.
  • introduction: A brief textual description providing an overview of the company.
  • business_scope: A textual description outlining the company’s main business operations.
  • employees: The number of employees working for the company.
@@ B.2 Real Stock Data Variables
  • main_business: A textual description of the company’s primary business activities.
@@
  • city: The city where the company’s headquarters is located.
  • name: The full name of the company.
  • industry: The industry to which the company belongs.
B.3 Empirical Analysis of Average Trading Volume for SSE 50 Constituents
Xueqiu and Guba are both prominent Chinese social media platforms catering to financial investors. Consequently, as illustrated in Table 4, we leverage transaction data from these two distinct sources:
  • Xueqiu: Data from Xueqiu covers the period from 2023-01-03 to 2023-12-06. The Xueqiu data corresponds to 639 real user accounts. These accounts are used to provide seed data, initializing user profiles and generating historical transaction records for the TwinMarket simulation.
  • Guba: Data from Guba covers the period from 2017-06-27 to 2024-06-03. The Guba data is used to inform stock recommendations within the TwinMarket environment.
@@ B.3 Empirical Analysis of Average Trading Volume for SSE 50 Constituents
Importantly, within the TwinMarket simulation, the Xueqiu data serves for user initialization, while the Guba data is utilized for stock recommendations. This difference in their roles, along with the different time periods, makes a direct comparison of the average trading volumes less meaningful. The primary purpose of presenting these statistics is to highlight the different scales and timeframes of the data used for different aspects of the simulation.
B.4 Information Source Details
@@
Announcements 5.6K 21283 15
@@
Table 5 and Figure 10 report the statistics of primary information source and the detailed top distribution over time in news.
Appendix C User Profile Initialization
C.1 Heterogeneous Investors in Financial Markets
Fundamentalist Traders.
The decision-making of rational traders typically follows classic financial theories, especially Modern Portfolio Theory (MPT) [23] and the Capital Asset Pricing Model (CAPM) [69]. Key factors considered by rational traders include expected returns and risk, that is, maximizing returns under an acceptable level of risk and minimizing risk under expected returns. According to MPT, traders also diversify risk by optimally allocating different assets. In our TwinMarket, these individuals are referred to as fundamentalist traders. We guide them step by step through prompts to analyze the fundamental indicators of assets, their corresponding valuations, expected returns, and the risks they are willing to tolerate.
Technical Traders.
In real-world markets, there is often information asymmetry or a lag in information processing, which leads some traders to base their trading decisions on historical data or price trends. These traders typically adopt a trend-following strategy, buying stocks or assets in an uptrend and selling those in a downtrend. Technical traders tend to trade very frequently and are more easily influenced by market sentiment and the behavior of other investors. In our TwinMarket, these individuals are referred to as technical traders. We guide them step by step through prompts to analyze technical trading signals in the market, as well as whether they should respond to market sentiment.
Behavioral Biases and Investor Heterogeneity.
In addition to differing trading strategies, expected returns, and risk preferences, financial markets are also significantly influenced by investors’ behavioral biases, which greatly contribute to market heterogeneity. According to behavioral finance theory, various psychological factors cause investors to make decisions that deviate from rational expectations, leading to inefficiencies and volatility in the market. In our TwinMarket, we have customized the following behavioral biases for each user:
  • Overconfidence Bias: The classical financial theory’s rational agent assumption predicts that markets would experience infrequent trading. However, in real markets, trading volume is substantial. Behavioral finance explains this excessive trading through overconfidence, where investors tend to overestimate their understanding of the market and their ability to predict market movements [70]. Even when they lack sufficient knowledge or information, they make decisions with excessive confidence.
  • Loss Aversion:Prospect theory suggests that people are twice as sensitive to losses as they are to gains [24]. As a result, investors are more likely to sell stocks that have appreciated in value too soon while holding on to stocks that have declined, a phenomenon known as the Disposition Effect [71].
  • Lottery Preference: Investors tend to favor investment options or decisions that resemble a lottery—those with low probability but high potential returns—while ignoring more stable options with lower returns [72].
  • Insufficient Diversification: Investors tend to prefer investments they are familiar with, such as domestic stocks or stocks of companies they know well, leading to suboptimal portfolio diversification [73].
C.2 Calculating Biases
@@
Leveraging the Xueqiu dateset encompassing 639 valid real-time trading users, and 11965 real-time trading records on the Chinese stock market, we employed methods similar to Sui’s pioneering work [26] to calculate behavioral biases related to individual trading activity.
Crucially, while both our study and Sui’s work utilized the Xueqiu platform and its trading data, our analysis is specifically restricted to a subset of stocks within the SSE 50 index, representing high-quality stocks.
Results, presented in Table 6, demonstrate similarities with those from the prior study (Table 7). Despite the more focused dataset, the observed behavioral patterns align closely, suggesting that the identified biases are robust across a range of stock selections on the Xueqiu platform. This finding demonstrates the scalability of the approach and lays a strong foundation for our subsequent construction of initial user information profiles.
C.3 Generating persona
@@
For each user, we included their gender, location, and social media follower count, as these dimensions have been shown to significantly influence investor behavior [70, 74]. Additionally, we incorporated four behavioral bias indicators that we previously calculated(Table 6). We sampled these user characteristics based on the joint distribution of the data in the Xueqiu dataset, thus generating the basic profiles for each user.
In the original dataset, the behavioral biases of real users were divided into three groups—“high,” “medium,” and “low”, to represent the severity of each bias. For biases with different severity levels, we created contextually realistic prompts to guide the user’s behavior accordingly. In Appendix E.1, we provide examples demonstrating how users exhibit human-like behavioral biases in their investment decisions.
For each generated user, we use the GPT-4o model to score their level of rationality. The top 40% of users, based on their rationality score, adopt a fundamental analysis investment strategy, while the remaining users follow a technical analysis strategy. Additionally, the top 10% of the most rational users are allocated ten times the initial funds compared to regular users. The distribution of strategy adoption and initial fund allocation is consistent with the findings from surveys of the A-share market [75].
C.4 Generating initial Transactions records
For each generated user, we find the most matching real user template and sample the historical trading data of the real user to generate the virtual user’s trading records. The purpose of this step is to initialize the user’s areas of interest (industries) and construct a User Graph based on the historical trading records.
C.5 Belief Initialization
To initialize the beliefs of users in a realistic manner, we first determine the mean belief for the market using the historical percentile of the BRAR sentiment index [^7] The BRAR sentiment index is a technical indicator derived from the balance of rising and falling stocks in the market, reflecting market sentiment. Data source: https://tushare.pro/. on the market’s starting date. We then sample belief values for each user across multiple dimensions using a Gaussian distribution centered around this mean. Finally, based on the sampled belief scores, we employ GPT-4 to generate contextually realistic and personalized belief content for each user, ensuring that their beliefs are both diverse and grounded in realistic market conditions.
Appendix D Framework Details
D.1 Dynamic Social Network
Figure 11: Chord diagram illustrating the user network, where each node represents a user, categorized by their preferred trading industry. Connections between users indicate behavioral similarity in trading patterns, highlighting clusters of users with aligned investment behaviors.
Figure 11 illustrates the evolution of user trading portfolios and their network connections over two time periods. Each node represents a user, categorized by their preferred trading industry, while connections indicate behavioral similarity in trading patterns. The shifting structure of the network highlights how user trading preferences evolve over time, leading to dynamic changes in market influence and information propagation.
This dynamic nature also demonstrates that users exhibit distinct preferences under different market conditions. For instance, in the network visualization of 2023-10-20, compared to 2023-06-20, there is a noticeable shift toward a stronger preference for the Consumer Goods and Technology & Communications industries. This change is likely driven by the influence of news releases, which may have contained a significant amount of positive information favoring these sectors. As a result, market sentiment shifted, prompting more users to adjust their holdings accordingly.
Furthermore, the structure of user connections adapts in response to these trading behavior changes. Users with similar trading patterns form tighter clusters, while those with diverging preferences become less connected. This reflects how information-driven market shifts lead to the reorganization of investor networks, where groups of traders dynamically form and dissolve based on evolving market narratives. Ultimately, this process underscores the crucial role of social influence and information diffusion in shaping financial market structures.
D.2 Post-propagation Chain
Mechanism.
Figure 12: Example of post-propagation chain.
When updating a post’s score, we consider both the breadth and depth of its propagation. If a user interacts with a root post by reposting or liking any of its propagated versions, this engagement contributes to the original post’s cumulative score. This ensures that highly engaging posts continue gaining visibility as they spread across the network.
Figure 12 illustrates an example of post propagation, where a single post can give rise to multiple propagation chains. Each branch represents a different path through which the information spreads, reinforcing the reach and influence of widely shared content.
Opinion Leader.
Figure 13: Example of post-propagation chains in social network
Under this mechanism, posts that resonate strongly with the broader community will achieve wider dissemination within the network. As users imitate behaviors and align with the perspectives of influential posters, their social connections become more centralized. This self-reinforcing process leads to the emergence of opinion leaders—users whose posts exert a disproportionately larger impact on the network.
As illustrated in Figure 13, high-degree centrality users gain influence as their posts receive increased engagement, leading to broader information diffusion. This effect is further supported by Figure 14, which shows that a few users experience significantly higher repost counts, reinforcing their role as key opinion leaders. Their ability to shape market sentiment and drive collective decision-making underscores the impact of social influence in financial markets, where information cascades can significantly alter trading behaviors and reshape network structures.
Figure 14: Trend of repost counts for all users in the 100-user network. The figure illustrates how the number of times users’ posts are reposted evolves over time, reflecting changes in information diffusion and engagement dynamics within the network. Notably, a few users exhibit significantly higher impact, indicating their role as key opinion leaders driving information propagation.
D.3 Belief

### v3

@@ E.1.1 SSE 50 and 10 Aggregated Index
In our simulation process, we mainly focus on the SSE 50 index [^6] More details about SSE 50 can be found in https://www.sse.com.cn/market/sseindex/overview/., a benchmark tracking the 50 largest, most liquid A-share stocks on the Shanghai Stock Exchange, reflecting the performance of China’s blue-chip market.
To provide a more granular view of sector-specific performance within the SSE 50, we introduce 10 aggregated indices. Each aggregated index, $I_{k}$, represents a distinct sector $k$ and comprises a subset of $N_{k}$ constituent stocks from the SSE 50. The value of each aggregated index is calculated as a weighted average of the price relatives of its constituent stocks. The weights are determined by each stock’s market capitalization at the start of the simulation ($t=0$) and remain fixed throughout the simulation.
Conceptually, the value of an aggregated index $I_{k}$ at simulation time $t$ is given by:
$I_{k,t}=I_{k,0}\cdot\sum_{j=1}^{N_{k}}\left(w_{j,0}^{(k)}\cdot\frac{P_{j,t}}{P_{j,0}}\right)$ (1)
where:
  • $I_{k,t}$ is the value of index $k$ at time $t$.
  • $I_{k,0}$ is the base value of index $k$ at $t=0$ (e.g., 100).
  • $w_{j,0}^{(k)}$ is the fixed weight of stock $j$ in index $k$, based on its market capitalization at $t=0$ relative to the total initial market capitalization of all stocks in index $k$.
  • $P_{j,t}$ is the price of stock $j$ at time $t$.
  • $P_{j,0}$ is the price of stock $j$ at $t=0$.
This method ensures the index reflects the performance of its components based on their initial market significance, with these significance levels (weights) held constant. The composition of each aggregated index is detailed in Table 10.
Table 10: Indices and Representative Stocks
Index Name Abbr. Count Representative Stock
Transportation and Logistics Index TLEI 2 Cosco Shipping Holdings
Manufacturing Index MEI 8 Longi Green Energy Technology
Chemical and Pharmaceutical Index CPEI 3 Jiangsu Hengrui Pharmaceuticals
Infrastructure and Engineering Index IEEI 3 China State Construction Engineering
Real Estate Index REEI 1 Poly Developments and Holdings
@@
Tourism and Service Index TSEI 1 China Tourism Group Duty Free
Consumer Goods Index CGEI 5 Kweichow Moutai
Technology and Telecommunications Index TTEI 10 Semiconductor Manufacturing International Corporation
Energy and Resources Index EREI 6 China Yangtze Power
Financial Services Index FSEI 11 Ping An Insurance
E.1.2 Stock Indicators
The simulation utilizes real-world stock data and company information for initialization and grounding. These variables, collected daily for each constituent stock of the SSE 50, are integral to various stages of the agent workflow and environment dynamics. They are categorized as follows:
General Indicators
These indicators provide basic identification and contextual information for each stock, primarily used for referencing and structuring information within the simulation, as detailed in Table 11.
Table 11: General Indicators
Indicator Description
stock_id The unique identifier code for each stock.
stock_name The full name of the company.
date The trading date associated with the record.
industry The industry classification derived from the aggregated indices (e.g., TLEI, MEI).
Fundamental Indicators
Primarily utilized by fundamentalist agents, these indicators reflect perceived company value based on financial performance and are key inputs for value-based analysis, primarily used for referencing and structuring information within the simulation, as shown in Table 12.
Table 12: Fundamental Indicators
Indicator Description
pe Price-to-Earnings Ratio: Compares the current stock price to the company’s earnings per share over the trailing twelve months. It reflects the market’s valuation of earnings.
pb Price-to-Book Ratio: Compares the current stock price to the company’s book value (net assets) per share. Used to assess valuation relative to the company’s accounting net worth.
ps Price-to-Sales Ratio: Compares the company’s market capitalization (or stock price) to its total revenue over the trailing twelve months.
dv Dividend Yield: Shows the annual dividend payout over the trailing twelve months as a percentage of the current stock price. Represents the dividend return component for shareholders.
Company Information Indicators
Providing static background details about the company’s structure, history, and operations, company information indicators are static during the simulation process, as detailed in Table 13.
Table 13: Company Information Indicators
Indicator Description
reg_capital The official registered capital of the company (RMB).
setup_date The date the company was legally established or incorporated.
introduction A brief narrative overview of the company’s business and market position.
business_scope A formal description outlining the range of business activities.
employees The reported number of individuals employed by the company.
@@ E.1.2 Stock Indicators
main_business A description highlighting the company’s primary revenue-generating activities.
@@
city The city where the company’s main headquarters or registered office is located.
Technical Indicators
Primarily utilized by technical agents, these indicators are derived from market trading activity (price and volume) and reflect trends within the market, as shown in Table 14.
Table 14: Technical Indicators
Indicator Description
close_price The final price recorded for the stock at the end of the trading day.
pre_close The closing price recorded for the stock on the immediately preceding trading day.
change The absolute difference between the current day’s close_price and the pre_close.
pct_chg The percentage change from the pre_close to the current day’s close_price.
vol The total number of shares traded during the trading day.
vol_5, vol_10, vol_30 The average daily trading volume over the preceding 5, 10, and 30 trading days, respectively. Used to gauge recent liquidity levels.
ma_5, ma_10, ma_30 The 5, 10, and 30-day simple moving averages of the closing price, adjusted backward for stocks. Used to identify price trends.
elg_amount_net The net value (buy minus sell) of large institutional orders executed during the day.
Data Handling Across Simulation Phases
Understanding the source and nature of these indicators before and during the simulation is crucial. Before the simulation begins (At initialization phase), all indicator data is populated using historical real-world data to ground the environment and agent profiles. However, once the simulation starts, the data sources and characteristics diverge significantly based on the indicator type, as summarized in Table 15.
Notably, fundamental valuation metrics are dynamically adjusted during the simulation to remain consistent with the emergent simulated stock prices to ensure fundamentalist agents can perform meaningful value analysis within the simulation’s context. The adjustment mechanism uses static base values (e.g., initial Earnings Per Share (EPS), initial Book Value Per Share (BVPS)) derived from real-world data at T=0 (Pre-simulation), combined with the current simulated close_price. Each day, the ratios are recalculated, for example:
Simulated $\text{Simulated }pb=\frac{\text{Current Simulated {Close\_Price}}}{\text{Initial BVPS}}$ (2)
This approach anchors simulated valuations to the initial real-world fundamentals while allowing the ratios to reflect the dynamic price discovery process within TwinMarket, enabling realistic agent decision-making based on perceived value.
Table 15: Comparison of Indicator Data Handling: Simulation Phase
Indicator Category Indicator Data Handling During Simulation
General Indicators Remain static references except the date advances according to the simulation’s internal clock.
Company Information Remain entirely static, providing fixed background context about the companies throughout the simulation run.
Fundamental Indicators Dynamically adjusted. Calculated by combining the current simulated close_price with the static fundamental base values loaded during initialization. For example, Simulated P/E = Simulated Price / Initial EPS.
Technical Indicators Entirely generated by the simulated Market Environment. Values like close_price and vol emerge directly from agent Trading actions processed via the order-driven matching mechanism. Derived indicators (MAs, average volumes) are calculated from this generated simulation data.
E.2 Transaction Details
TwinMarket leverages real-world transaction data from two prominent Chinese social media platforms for financial investors: Xueqiu and Guba. These platforms provide valuable insights into retail investor behavior and market discussions, serving distinct roles in our simulation framework.
Xueqiu Data for User Profile Initialization
Transaction records and associated user information from 639 real Xueqiu user accounts, covering the period from 2023-01-03 to 2023-12-06, serve as the primary dataset for initializing agent profiles in TwinMarket. This choice is strategically aligned with established academic research[27], which utilized Xueqiu data for behavioral analyses. By grounding our bias calculation (detailed in Section F.2) on the same data source and a comparable analytical approach, we ensure that the behavioral characteristics embedded in our agents are consistent with peer-reviewed findings. This dataset is crucial for the initialization phase of TwinMarket, as detailed in Section F.
Guba Data for Recommendation System Training
Transaction data from Guba, which spans a more extensive period from 2017-06-27 to 2024-06-03, provides a broader view of trading activities. Within TwinMarket, this larger dataset is primarily used to inform and train the stock recommendation system. The patterns of stock popularity and trading frequency observed in the Guba data help simulate how certain stocks might become "hot" or be recommended to users in a dynamic information environment.
Data Summary and Context
Table 16 summarizes the key characteristics of the data sourced from these platforms.
@@ E.2 Transaction Details
E.3 Information Source Details
To provide a realistic information environment for the agents within TwinMarket, we incorporate two primary types of external information: News Articles and Company Announcements. These sources simulate the flow of public information that influences investor sentiment and decision-making.
Table 17 presents a quantitative overview of these information sources. News articles, with over a million instances and an average of nearly 2,900 articles per day, represent the broad and continuous stream of general market and economic information. Company announcements, while less frequent (averaging 15 per day), are typically much longer and more specific, often containing material information directly relevant to individual stocks.
@@ E.3 Information Source Details
Company Announcements 5.6K 21283 15
To further illustrate the temporal distribution and thematic composition of the news articles, Figure 25 displays the average number of news items aggregated over 14-day intervals throughout 2023. The stacked bars break down the news into key categories such as society, economic, entertainment, and other. This detailed breakdown of information sources ensures that agents in TwinMarket are exposed to a rich and varied stream of information, mimicking the complex data environment faced by real-world investors.
@@
Appendix F Technical Details II: User Profile Initialization
The creation of realistic and diverse agent profiles is fundamental to the TwinMarket simulation. This section details the multi-stage process of initializing these user profiles, grounding them in real-world investor characteristics, established financial theories, observed behavioral biases, and initial market beliefs. Our goal is to populate the simulation with agents whose initial states reflect the heterogeneity found in actual financial markets.
F.1 Foundations: Heterogeneous Investors in Financial Markets
To realistically simulate investor behavior, TwinMarket incorporates heterogeneity by modeling distinct investor archetypes. Our primary focus is on capturing a spectrum of investment approaches commonly observed in financial markets. While institutional investors play a significant role globally, our simulation is specifically contextualized within China’s A-share market. This market presents a unique characteristic where institutional investors often exhibit trading patterns and behavioral tendencies more akin to those of retail investors [75, 76].
This market-specific consideration informs our modeling strategy. While we primarily differentiate agents based on two dominant analytical styles discussed in financial literature – Fundamental Traders and Technical Traders, many of the behavioral biases and social interaction patterns simulated are broadly applicable due to this observed blurring of roles. Furthermore, to account for the varying impact of different market participants, we approximate the influence of larger players or more capitalized entities through agents initialized with significantly higher capital (as detailed in Appendix F). This approach allows us to observe differential market impact while remaining consistent with the A-share market’s particular participatory dynamics.
Fundamental Trader
These traders base their investment decisions on an analysis of a company’s intrinsic value, typically derived from its financial health, earnings potential, and broader economic conditions. Their approach aligns with classic financial theories like Modern Portfolio Theory (MPT) [24] and the Capital Asset Pricing Model (CAPM) [77], emphasizing risk-return optimization and diversification. In our TwinMarket framework, fundamentalist agents are guided through prompts to analyze fundamental indicators (see Appendix E.1.2), assess valuations, and consider expected returns and risks.
Technical Traders
These traders focus on historical price patterns, trading volumes, and market trends to predict future price movements, often operating under the assumption that market psychology and past trading activity can indicate future performance. They tend to engage in more frequent trading and may be more susceptible to short-term market sentiment. In our TwinMarket framework, technical traders are prompted to analyze technical signals (see Appendix E.1.2) and react to perceived market momentum.
F.2 System Prompt Generation
To ensure the behavioral realism of our agents, we ground their initial psychological biases in empirical data. This process involves quantifying specific behavioral tendencies using transaction histories from real investors.
Data Acquisition
Leveraging the Xueqiu dateset encompassing 639 valid real-time trading users, and 11965 real-time trading records on the Chinese stock market(see Appendix E.2). Crucially, while both our study and Sui’s work utilized the Xueqiu platform and its trading data, our analysis is specifically restricted to a subset of stocks within the SSE 50 index, representing high-quality stocks.
Incorporating Behavioral Biases
Beyond rational decision-making frameworks, real-world investor behavior is significantly influenced by psychological biases. Following Sui’s work which also utilized transaction records and associated user information from Xueqiu[27], we incorporate several well-documented biases into our agent profiles to contribute to the initialization of agent’s unique decision-making profile:
  • Disposition Effect: The tendency for investors to sell assets that have appreciated in value too early, while holding onto assets that have depreciated for too long, often in the hope of a recovery [78].
  • Lottery Preference: An inclination towards investments that offer a small chance of a very large payoff, similar to a lottery ticket, even if such investments have a negative expected value or are otherwise sub-optimal from a risk-return perspective [79, 80].
  • Underdiversification: The practice of holding a portfolio that is not sufficiently diversified across different assets or asset classes, often due to familiarity bias (preferring known investments), thereby exposing the investor to unnecessary idiosyncratic risk [81].
  • Turnover: Refers to the frequency with which an investor buys and sells assets in their portfolio. High turnover can be indicative of overconfidence, excessive reaction to noise, or attempts at market timing, and often leads to higher transaction costs [82].
Validation of Sample Representativeness
Given our sample size and acquisition method, a crucial step is to validate whether our dataset, despite its relatively small scale, can yield meaningful behavioral insights. We achieve this by comparing the descriptive statistics of the calculated biases from our 639-user Xueqiu sample with those reported in the larger-scale study by Sui[27].
Table 18 presents the summary statistics for these biases from our sample, while Table 19 shows the reference statistics from Sui. We can conclude from the result that the behavioral patterns observed in our smaller sample exhibit notable similarities in these investor biases compared to the findings from the larger reference study. This congruence in the statistical profiles of key behavioral biases suggests that our independently collected Xueqiu dataset, even with its more limited scale, captures fundamental behavioral tendencies that are consistent with those identified in broader, more extensively resourced studies. This finding supports the appropriateness of using our data and analytical approach for initializing the behavioral parameters of the agents in TwinMarket.
@@ F.2 System Prompt Generation
Generation of Agent Personas
With the foundational investor types defined and behavioral biases quantified, we proceed to generate individual agent personas:
  1. 1.
@@
    Demographic and Social Attributes: For each user, we included their gender, location, and social media follower count, as these dimensions have been shown to significantly influence investor behavior [83, 84].
  2. 2.
    Assigning Behavioral Bias Levels: Each agent is assigned a level (e.g., “high”, “medium”, “low”) for each of the four calculated behavioral biases based on their quantified score from the Xueqiu data. These levels inform the textual prompts used to guide the LLM in embodying specific biases. An example of a generated persona incorporating these elements is shown in Figure 26.
  3. 3.
    Investment Strategy Assignment: We use GPT-4o to assess the overall “rationality” of each agent based on their combined demographic and bias characteristics. The top 40% of agents deemed more “rational” are assigned a fundamental analysis strategy, while the remaining 60% adopt a technical analysis strategy.
  4. 4.
    Initial Capital Allocation: To reflect potential differences in sophistication or risk tolerance associated with rationality, the top 10% of the most “rational” agents are allocated ten times the initial capital compared to other agents. The proportions used for both strategy assignment (as described in point 3) and initial capital allocation are consistent with observed patterns in A-share market surveys [75].
F.3 Initialization of Trading Records and Beliefs
After establishing agent personas, strategies, and behavioral biases, we initialize their starting transaction histories and market beliefs to ensure a comprehensive and realistic initial state for each agent.
Generating Initial Transactions Records
To provide each agent with a plausible starting point for their investment journey and social interactions, we perform the following:
  1. 1.
    Synthesis of Initial Transaction Records: To equip each pre-defined synthetic agent with a plausible investment starting point, we generate a sequence of initial transaction records.
    • Firstly, trading patterns are extracted from an anonymized dataset of real-world investor transactions, which notably includes industry-specific details for each trade. This analysis serves to capture key characteristics for each real investor profile, such as their typical trade frequency, preferences for specific industries (modeled as a probability distribution), common trading directions (e.g., buy or sell tendencies), and typical trade volume statistics.
    • For each synthetic agent, whose detailed persona and guiding prompt are already established, a trading pattern from a randomly selected real investor profile is chosen to serve as a behavioral template.
    • A new sequence of synthetic transactions is then algorithmically generated for the agent, designed to stochastically mirror the chosen template’s characteristics. Specifically, the number of generated trades is based on the template’s historical trade count, subject to minor random variations. Trade dates are resampled to fall within the historical period covered by the real data, and crucial transaction details—such as the industry sector traded, the direction of the trade, and its volume—are determined by sampling from the template investor’s observed preferences and volume distributions.
    • These generated transaction records, forming the initial trading history for each agent, are then compiled. This process effectively initializes each agent’s notional portfolio holdings and establishes observable early investment leanings, such as preferred industries or typical trading frequency.
  2. 2.
    Foundation for Initial Social Network Seeding: The synthesized initial transaction histories serve as the primary foundation for seeding the initial social network. Specifically, similarities derived directly from these transactions, such as shared investment interests in particular industries (identified from the industry field in the generated transaction data), form the basis for establishing these initial connections. This transaction-derived data is then utilized to construct an initial social network graph. Agents demonstrating overlapping investment foci (e.g., frequent trading in the same sectors) might be assigned a higher initial probability of connection or mutual influence. This approach aims to simulate pre-existing communities of interest or implicit relationships based purely on observed market activity within the generated transaction histories.
Belief Initialization
An agent’s initial beliefs about the market are critical determinants of their early actions within the simulation. We initialize these beliefs through a structured, data-grounded, and LLM-enhanced process:
  1. 1.
    Market Sentiment Anchor Determination: We first establish a baseline for the overall market sentiment. This is achieved by taking the historical percentile, denoted as $P\in[0,1]$, of the BRAR sentiment index corresponding to the simulation’s designated start date. This percentile $P$ is then mapped to an initial mean market sentiment score, $\mu_{\text{market\_sentiment}}$, on a 0-to-10 scale. This ensures that the 50th percentile ($P=0.5$, representing the historical average sentiment) corresponds to a neutral score of 5. For instance, if the BRAR value on the simulation’s start date corresponds to its 70th historical percentile ($P=0.7$), this yields $\mu_{\text{market\_sentiment}}=7$.
  2. 2.
    Individual Belief Score Sampling: For each agent $i$, belief scores on a 0-to-10 scale across multiple predefined dimensions (e.g., outlook on economic fundamentals, market valuation levels, short-term market trends, sentiment of surrounding investors, and self-assessment of investment skill, see Appendix G.1 for a detailed list of belief dimensions) are then independently sampled from a Gaussian distribution. The mean of this Gaussian distribution for each dimension is set to the $\mu_{\text{market\_sentiment}}$ derived in the previous step. The variance for this sampling is proxied by the variance of the BRAR index itself, calculated over a 20-trading-day window centered around the simulation’s start date.
  3. 3.
    LLM-Powered Belief Narrative Generation: The sampled numerical belief scores (on a 0-to-10 scale for each dimension), along with the agent’s established persona (including their demographics, assigned behavioral biases, and investment strategy), are then fed as structured input into a GPT-4o prompt. This prompt is designed to interpret these 0-to-10 scores, with a score of 5 representing a neutral stance for each respective dimension, thereby guiding the LLM to generate a coherent, contextually rich, and personalized textual narrative articulating the agent’s initial beliefs.
This comprehensive initialization process shown in Figure 27 aims to create a diverse population of agents with plausible initial characteristics, strategies, biases, trading histories, and beliefs, thereby setting a rich stage for the observation of emergent socio-economic dynamics within TwinMarket.
Figure 27: Pipeline of User Profile Initialization.
Appendix G Technical Details III: TwinMarket Framework Mechanism
G.1 Agent Belief Structure and Representation